Find the solution of these problems diagrammatically and algorithmically.
1. 1/5+4/5.
2. 1/5+3/10.
3. 1/3+1/2.
4. 1/3+2/5.
5. 1/4+5/6.
1
Expert's answer
2020-09-29T18:53:00-0400

1) 1/5+4/5= 5/5=1


2)1/5+3/10

2/2×1/5 +3/10

=2/10+3/10

=5/10=1/2


3)1/3+1/2

=2/2×1/3+3/3×1/2

=2/6+3/6=5/6


4) 1/3+2/5

=5/5×1/3 +3/3×2/5

=5/15+6/15

=11/15


5)1/4+5/6

=3/3×1/4+2/2×5/6

=3/12+10/12

=13/12